On January 1, Hemi Corporation issued $350,000 of 9-year, 8% bonds for $309,086 yielding an effective interest rate (yield to maturity) of 10%. Interest is paid-semi-annually on June 30 and December 31.

Required:

1) Show computations to confirm the issue price

2) compute an amortization schedule for the 9 years of the bonds' life.

3) Give journal entries for: (A) bond issuance, (B) the semi-annual interest payment on June 30 of the first year, and (c) the semi0annual interest payment on December 31 of the first year.
Answer #1
Bond characterstics Amount
1) Principal 350,000
interest      14,000
Market interest rate 5%
periods to maturity 18
issue price 309,086
principal * PV of $1 at 5% for 18 yrs =
350,000 * 0.41552        = 145432
interest * PV of ordinary annuity at 5%=
14000 * 11.68959 = 163654
bond issue price 309086
2) Amortization Schedule
Date cash intt inttt exp Discount Carrying
amortized value
1/1/ 309,086
30-Jun      14,000 15454 1454 310541
31-Dec      14,000 15527 1527 312068
30-Jun      14,000 15603 1603 313671
31-Dec      14,000 15684 1684 315355
30-Jun      14,000 15768 1768 317122
31-Dec      14,000 15856 1856 318978
30-Jun      14,000 15949 1949 320927
31-Dec      14,000 16046 2046 322974
30-Jun      14,000 16149 2149 325122
31-Dec      14,000 16256 2256 327378
30-Jun      14,000 16369 2369 329747
31-Dec      14,000 16487 2487 332235
30-Jun      14,000 16612 2612 334846
31-Dec      14,000 16742 2742 337589
30-Jun      14,000 16879 2879 340468
31-Dec      14,000 17023 3023 343492
30-Jun      14,000 17175 3175 346666
31-Dec      14,000 17333 3333 350000
3) Date Account titles & Explanations Debit Credit
1-Jan Cash 309,086
discount on bonds payable 40,913
Bonds payable 350,000
30-Jun interest expense 15454
Discount on bonds payable 1454
cash      14,000
31-Dec interest expense 15527
Discount on bonds payable 1527
cash      14,000